### v1

# Observations of high-frequency spectral peaks from in-situ waves in ice data: evidence for nonlinear waves in ice triad interactions?

Observations of high-frequency spectral peaks from in-situ waves in ice data: evidence for nonlinear waves in ice triad interactions?
Jean Rabault
Norwegian Meteorological Institute
Oslo, Norway
jean.rblt@proton.me
&Joey Voermans
The University of Melbourne
Melbourne, Australia
jvoermans@unimelb.edu.au
&Takehiko Nose
The University of Tokyo
Kashiwa, Japan
tak.nose@edu.k.u-tokyo.ac.jp
&Graig Sutherland
Environment Climate Change Canada
Dorval, Canada
graig.sutherland@gmail.com
&Alexander Babanin
The University of Melbourne
Melbourne, Australia
a.babanin@unimelb.edu.au
&Takuji Waseda
The University of Tokyo
Kashiwa, Japan
waseda@edu.k.u-tokyo.ac.jp
&Tsubasa Kodaira
The University of Tokyo
Kashiwa, Japan
kodaira@edu.k.u-tokyo.ac.jp
&Atle Jensen
University of Oslo
Oslo, Norway
atlej@math.uio.no
&Lars Willas Dreyer
University of Oslo
Oslo, Norway
larswd@math.uio.no
&Øyvind Breivik
Norwegian Meteorological Institute
Bergen, Norway
oyvindb@met.no
&Gaute Hope
Norwegian Meteorological Institute
Bergen, Norway
gauteh@met.no
&Malte Müller
Norwegian Meteorological Institute
Oslo, Norway
maltem@met.no
&Zhaohui Cheng
Uppsala University
Uppsala, Sweden
zhaohui.cheng@geo.uu.se
&Lichuan Wu
Uppsala University
Uppsala, Sweden
lichuan.wu@geo.uu.se
&Aleksey Marchenko
The University Centre in Svalbard
Svalbard, Spitsbergen, Norway
alekseym@unis.no
&Petra Heil
University of Tasmania
Hobart, Australia
petra.heil@utas.edu.au &Karsten Trulsen
University of Oslo
Oslo, Norway
karstent@math.uio.no
Abstract
The propagation of waves through the marginal ice zone (MIZ) and deeper into pack ice is a key phenomenon that influences the breakup and drift of sea ice. Waves in ice propagation can be characterized by the associated dispersion relation, which describes both the speed and wavelength of the waves and their attenuation. When waves in ice propagate through a solid, non-cracked, thick enough sea ice cover, significant flexural elastic effects can be present in the dispersion relation. This results in a dispersion relation that opens up for 3-wave interactions, also known as wave triads.
Here, we report the observation of high-frequency spectral peaks in the power spectral density of waves in ice spectra. We show, in two timeseries datasets, that the presence of these high-frequency peaks is accompanied by high values for the spectral bicoherence. This is a signature that the high-frequency peak is phase-locked with frequency components in the main spectral energy peak, and a necessary condition for nonlinear coupling to take place. Moreover, we show for a timeseries dataset that includes several closely located sensors that the dispersion relation recovered from a cross-spectrum analysis is compatible with the possible existence of wave triads at the same frequencies for which the bicoherence peak is observed. In addition to these observations in timeseries datasets, we show that similar high-frequency peaks are observed from additional, independent datasets of waves in ice power spectrum densities transmitted over iridium from autonomous buoys.
These results suggest that nonlinear energy transfers between wave in ice spectral components are likely to occur in some waves and sea ice conditions. This may enable redistribution of energy from weakly damped low-frequency waves to more strongly attenuated higher-frequency spectral components, which can contribute to energy dissipation in the ice. However, more data are needed to offer definite conclusions about the practical importance of this effect in real-world conditions. We suggest several in-situ measurements, numerical investigations, and laboratory experiments to further investigate these phenomena.
1 Introduction
The Marginal Ice Zone (MIZ) is the area of sea ice that is strongly influenced by the propagation of open ocean waves. These waves can, in particular, lead to breakup of sea ice collins2015situ ; voermans2020experimental , which can dramatically alter sea ice conditions and greatly influences sea ice drift sutherland2022estimating ; thomson2022wave and melting thomson2014swell ; roach2019advances .
The propagation of waves in ice-covered areas is characterized by the wave dispersion relation, which in the absence of currents is formulated as an equation $D(k,f)=0$, where $k$ is the wave number modulus (in rad/m), $f$ is the wave frequency (in Hz; the wave angular frequency $\omega=2\pi f$ in rad.Hz is often used in place of $f$, and vice versa), and $D$ is the equation of the dispersion relation. In addition, parameters such as the water depth $H$ (in m), water volumetric mass density $\rho_{w}$ (in kg/m3), and the properties of the sea ice (in particular, the sea ice Young modulus $E$ in Pa, Poisson coefficient $\nu$, compressive stress $P$ in Pa, ice thickness $h$ in m, sea ice volumetric mass $\rho_{i}$ in kg/m3), play a role in the dispersion relation. More physical parameters may enter the dispersion relation if the influence of additional physics is taken into account under specific conditions. The wave number is, in general, a vector $\mathbb{k}(f)=\mathbb{k}_{r}(f)+i\mathbb{k}_{i}(f)$, where the real part $\mathbb{k}_{r}(f)$ characterizes the relation between the wavelength and the wave frequency, from which the phase and the group speed can be derived, while the imaginary part $\mathbb{k}_{i}(f)$ characterizes the wave attenuation.
Numerous papers have discussed the dispersion relation for waves in ice; see, e.g. squire1995ocean ; voermans2021wave . Often, the real (propagation properties) and imaginary (attenuation) parts of the dispersion relation are considered separately from each other, in which case $k=|\mathbb{k}_{r}|$ stands for the real part of the wavenumber, and one may write $\mathbb{k}$ instead of $\mathbb{k}_{r}$ as a shorthand notation. For example, greenhill1886wave wrote down (though it may have been known even earlier) a now classical form for the real part of the wave dispersion relation in ice under the assumption that the ice acts as a thin elastic sheet where flexural stiffness effects are present:
$\omega^{2}=\frac{gk+Bk^{5}-Qk^{3}}{\coth(kH)+kM},$ (1)
where $B=\frac{Eh^{3}}{12\rho_{w}(1-\nu^{2})}$, $Q=\frac{Ph}{\rho_{w}}$, $M=\frac{\rho_{i}h}{\rho_{w}}$.
In the following, we will consider the case where there is no significant contribution arising from the compressive stress, and the wave frequency is low enough that the mass loading effect is negligible with respect to other terms, so that the dispersion relation reduces to:
$\omega^{2}=\frac{gk+Bk^{5}}{\coth(kH)}.$ (2)
Interestingly, the term representing the effect of elasticity (i.e. $Bk^{5}$) makes it easy for the nonlinear resonant wave triad necessary conditions between 3 wave components (1,2,3)(1,2,3)( 1 , 2 , 3 ), i.e. (using the shorthand notation mentioned above and omitting the r indexes for simplicity):
$\begin{cases}f_{1}+f_{2}&=f_{3}\\ \mathbb{k}_{1}+\mathbb{k}_{2}&=\mathbb{k}_{3},\end{cases}$ (3)
to be fulfilled phillips1966dynamics ; phillips1981wave ; madsen1993bound , as we will discuss later in this manuscript. Although triads and other resonant interactions have, to the best of our knowledge, not been discussed extensively on the basis of field measurements, a few theoretical works can be found that discuss their possible existence marchenko1999stability ; marchenko1999stability2 . The only reported observations of high-frequency spectral peaks we know of, and that may be similar to ours, are the ones from johnson2021observing , though these seem to match a 4th order, rather than 2nd order (as we observe here, see below) high-frequency spectral peak, and it is uncertain what their origin is and whether they may come from the effect of wind per se, or can be the result of some nonlinear wave energy redistribution.
Regarding wave attenuation, the most widely accepted consensus, based on theoretical arguments, laboratory experiments, and empirical measurements in the ice wadhams1973attenuation ; yu2022 ; squire2020ocean ; loken2021wave ; zhao2015modeling ; zhao2015wave , is that the amount of wave energy dissipation at each frequency is proportional to the local amount of wave energy. This leads to an exponential attenuation of the waves as they propagate in the sea ice and is consistent with the wavenumber formalism introduced above, resulting in:
$a(f,\mathbb{x},t)=a(f,0,t)|e^{i(\mathbb{k}(f)\cdot\mathbb{x}-\omega t)}|=a(f,0,t)e^{-k_{i}(f)x_{\mathrm{eff}}},$ (4)
where $t$ indicates time (which disappears under a stationary regime assumption), and $a(f,\mathbb{x},t)$ is the amplitude of the wave component with frequency $f$ at a location $\mathbb{x}$ and time $t$, corresponding to an effective propagation depth $|\mathbb{k}_{i}(f)\cdot\mathbb{x}|=k_{i}(f)x_{\mathrm{eff}}$ in the sea ice.
Generally, $k_{i}(f)$ is an increasing function of wave frequency $f$ meylan2014situ ; yu2022 . Currently, there is general agreement in the literature on the existence of a scaling $k_{i}(f)\propto f^{n}$, with the value of the exponent $n$ reported to be typically around 4.5 yu2022 , although there is some variability in the literature depending on the exact choices for the wave in ice damping parameterization and the dataset(s) over which this relation is fitted.
However, there is still a debate in the literature on what physical mechanism, or combination of mechanisms, physically determines the attenuation coefficient $k_{i}(f)$, with effects such as scattering bennetts2010three ; kohout2008elastic ; montiel2016attenuation ; zhao2016diffusion ; bennetts2024thin , viscous dissipation zhao2015modeling ; sutherland2019two ; rabault2017measurements ; marchenko2018wave , turbulence voermans2019wave ; smith2020pancake , floe-floe interaction herman2019wave ; loken2022experiments ; herman2019wave ; smith2020pancake , viscoelasticity in ice mosig2015comparison ; zhang2021theoretical ; zhao2018three ; marchenko2021laboratory , and overwash pitt2022model ; zha2023numerical , regularly discussed. Although several formulations have been proposed and implemented to parameterize waves in ice damping (for example, at least 10 different parameterizations are available in WaveWatch III, tolman1991third ; rogers2013implementation ; KGulla ), and progress has been reported recently using fitted parameterizations based on scaling laws to unify data sets that span several orders of magnitude of waves in ice parameters yu2022 , models are still largely imperfect at representing waves in ice and, in particular, the wave in ice attenuation $k_{i}(f)$ KGulla ; rabault2024situ . Possibly, different combinations of dissipation mechanisms may play a role in different wave and ice conditions, making the problem complex and difficult to disentangle and solve in a general way herman2024apparent . If triads exist for waves in ice, this may add an additional possible dissipation mechanism by stealing energy from the main lower-frequency spectral components that are weakly attenuated and moving it into higher frequency components where energy dissipation is stronger.
The ongoing discussions and uncertainties about the details of the mechanisms causing the wave attenuation and the exact formula to use highlight the need for more ground-truth direct in-situ observations to help constrain and calibrate theoretical models. This has been the focus of the waves in ice community for many decades wadhams1973attenuation ; doble2006wave ; thomson2013sea ; kohout2016situ , and our groups have recently contributed significant amounts of such data rabault2023dataset ; rabault2024position ; muller2025distributed ; rabault2022tracking using new generations of open source, low-cost instruments based on the OpenMetBuoy (OMB) and closely related line of designs rabault2016measurements ; rabault2017measurements ; rabault2020open ; marchenko2019wave ; loken2021wave ; rabault2022openmetbuoy ; kodaira2024affordable ; kodaira2022development ; hope2025sfy ; MoreRoomattheTopHowSmallBuoysHelpRevealtheDetailedDynamicsoftheAirSeaInterface . The availability of larger in-situ datasets allows, in turn, to perform more detailed and representative analysis of real-world ground-truth data, which has contributed to reveal interesting dynamics dreyer2024direct ; rabault2024situ ; nose2024observation ; nose2023comparison .
In the present work, we study several diverse wave in ice datasets that we have gathered over the past ten years. For 2 of these datasets, we were able to obtain raw timeseries of the ice motion at 10Hz. For the other datasets, only the wave power spectrum density (PSD) transmitted over the Iridium satellite communication network is available. We show that in all of these datasets we can, now and then, observe clear high-frequency peaks in the PSD of the wave in ice elevation. In the case when timeseries are available, we demonstrate that these high-frequency peaks are accompanied by peaks in the bicoherence, and that they are compatible with the effect of nonlinear wave energy transfer through wave triads. This potentially opens for a novel nonlinear energy transfer and ultimately dissipation mechanism, as pointed out above.
The organization of the manuscript is as follows. In section 2, we present the datasets used and the methodology applied to analyze these. In section, 3, we present the results of our analysis. In section 4, we discuss our findings and their impacts. Finally, we offer some conclusions and suggest ways to extend our present investigation.
2 Data and methods
In this section, we present both the data we use and the methodologies applied to process these data.
2.1 Deployments considered
The data used in the following come from several deployments in sea ice under a variety of conditions. Two different categories of data were obtained: in two datasets, we were able to recover the instruments, so full timeseries are obtained. However, in most deployments the instruments could not be recovered and only the wave spectra transmitted by the buoys are available. The datasets are described in more detail in the following.
2.1.1 Timeseries data: Tempelfjorden 2015 (T15)
The first and most comprehensive dataset comes from the 2015 waves in ice measurements performed on landfast ice in Tempelfjorden, Svalbard, using a MOXA logger sutherland2016observations ; rabault2016measurements . A total of 5 Vectornav VN100 IMU (Inertial Motion Unit)-based sensors were deployed. Of these, the data from three locations have proved to work well from a technical point of view and produce useful data without measurement dropouts: these correspond to the IMUs with IDs VN1005, VN1002, and VN1007 (from closest to the open water to deepest into the ice), which we will consider in the following.
The VN100 data acquired during this deployment contain the raw X, Y and Z (that is, in the reference frame of the VN100 IMUs) sea ice motion data. In addition, the data are obtained by point-wise sampling of the internal readings of individual sensors at 10 Hz, while the sensors themselves run internally at 800 Hz. This means that we need to apply Kalman filtering a posteriori to the data sutherland2017method (since the sensors frame of reference is used) and that the noise level is slightly higher than in other deployments (since sampling rather than averaging is used). The data and processing are openly available; see the details in Appendix A.
This dataset provides time-synchronized measurements from 3 sensors at 10Hz. The first sensor (ID VN1005) is around 100 m from the ice edge. The second sensor (ID VN1002) is around 65 m farther in the ice than ID VN1005 along the wave propagation direction. The third sensor (ID VN1007) is around 7 m farther in the ice than ID VN1002 along the wave propagation direction. This dataset makes it possible to perform both bicoherence and cross-spectrum analyses between sensors, as will be discussed below.
The timeseries considered here continuously cover a period of around 14.5 hours. When analyzing these timeseries, the signals are split into segments of length 214 points to make Fast Fourier Transform (FFT) operations fast. This corresponds to around 27 minutes of data per segment. In the following, we will refer to the segment numbers when performing the analysis.
2.1.2 Timeseries data: Yermak plateau 2020 (Y20)
The second dataset comes from the OMB-v2018 deployment performed in 2020 over the Yermak plateau from the hovercraft R/H Sabvabaa rabault2023dataset . There, a total of six OMB-v2018s, each containing a VN100 IMU, were deployed on sea ice. Later, one of the buoys, with ID 18954, was found and recovered after being stranded on the northwest coast of Iceland dreyer2024direct . Therefore, the raw timeseries of the VN100 motion data recorded by this instrument are available from the on-board SD card. The IMU configuration is improved compared to the T15 dataset. More specifically, these VN100 IMU data contain Kalman-filtered (by the processor on board of the IMU package) 10 Hz averaged (by performing an 80-point averaging, instead of individual sample extraction, which provides lower noise) acceleration data in the Earth frame of reference rabault2020open ; dreyer2024direct . The data are openly available; see the details in Appendix A.
Timeseries are available each time the instrument wakes up to perform a wave measurement, that is, around every 4 hours. Each timeseries is around 21 minutes long and is referred to as a data segment, similar to the T15 deployment. In the following, we will refer to the segment numbers (the numbering is based on the file numbers on the recovered SD card, see the raw data release) when performing the analysis.
2.1.3 Spectra over Iridium
The third dataset includes data from Sofar Spotter raghukumar2019directional , OMB-v2018 rabault2020open , and OMB-v2021 rabault2022openmetbuoy deployments, which collected data in both the Arctic and Antarctic. In these deployments, the buoys could not be recovered, so only the 1-dimensional power spectral density (PSD) transmitted over iridium satellite communications is available.
More specifically, data from the following deployments are used:
  • The Davis 2020 (D20) Antarctic deployment rabault2023dataset . In this deployment, two OMB-v2018 and two Sofar Spotter were deployed on Antarctic landfast ice along a transect perpendicular to the broken ice edge, close to the Davis Research Station.
  • The Casey 2022 (C22) Antarctic deployment. In this deployment, the OMB-v2021s were deployed on Antarctic landfast ice, close to the Casey Research Station.
  • The Yermak 2020 Arctic deployment. This is the same deployment as Y20 discussed above rabault2023dataset , but data from other buoys are considered.
  • The 2022 CIRFA-UIT Arctic deployment. This deployment was performed in the Western Fram straight, West of Greenland. 19 OMB-v2021s were deployed from the R/V Kronprins Haakon. The instruments were deployed in April and May 2022, and data transmissions took place until December 2022 rabault2024position .
  • The 2022 AWI-UTOKYO Arctic deployment. This deployment was performed with 15 OMB-v2021s. The OMBs were deployed on the sea ice northwest of Svalbard in July 2022 and transmitted until October 2022 rabault2024position .
The goal of considering so many different deployments is to illustrate that high-frequency peaks in the waves PSD can be reliably observed in a variety of locations (covering both the Arctic and Antarctic), sea ice conditions (drift ice and landfast ice of various thicknesses, age, and properties), and seasons, and are not just a particularly seldom curiosity. Moreover, three different types of waves in ice instrument have been used to collect these data, relying on different measurement techniques and different signal processing implementations. In particular, while the Sofar Spotter is a close source design, the Spotter version used to generate the data in the Davis 2020 deployment relies, to the best of our knowledge, on GPS-based Doppler velocity measurements, while the OMBs rely on IMU-based measurements (and the OMB-v2018 and OMB-v2021 use different actual physical sensors, and different implementations of the signal processing routines). As a consequence, given the diversity of instruments, actual hardware sensors and measurement methods, and signal processing implementations used, consistent observation of specific dynamics across these instruments and deployments is very unlikely to be the result of systematic sensors or processing artifacts.
2.2 Analysis of raw timeseries
In this subsection, we describe the methodology used to analyze the raw timeseries data available from the T15 and Y20 deployments.
Note that, since all signals considered here are single real-world realizations of an underlying stochastic process (assumed to be stationary within each data segment considered), the Fast Fourier Transform- (FFT-) based formula that we use are strictly speaking only (stochastic) estimators for the underlying "true" characteristics of these signals, which are formally defined based on the underlying (and hidden) autocorrelation and cross-correlation functions (see classical textbook-like introductions to the topic, such as https://en.wikipedia.org/wiki/Autocorrelation and https://en.wikipedia.org/wiki/Wiener%E2%80%93Khinchin_theorem, both accessed 2025-07). However, for simplicity of writing and similar to commonly used conventions in this field, we refer to the FFT-based estimator formulae as the quantities, rather than the estimators for the quantities, that are considered.
2.2.1 Kalman processing of the T15 data
For the T15 data, Kalman processing must be applied to the raw IMU data measured in the IMU reference frame to obtain the ice motion in the Earth (North, East, Down) reference frame. This is done using the same open source Kalman filter that is used on the OMB-v2021 rabault2022openmetbuoy . The Kalman filter itself is available at https://github.com/adafruit/Adafruit_AHRS and https://github.com/gauteh/ahrs-fusion (both accessed 2025-07), and all code and instructions are available along with the data released with this manuscript; see Appendix A.
2.2.2 Power Spectral Density calculation
The wave Power Spectral Density (PSD) is calculated from the downward acceleration component in the Earth reference frame. When computing the PSD, we use the Welch method welch1967use , typically with an FFT length of 2048 points, 50% overlap, and a Hann windowing function. This is a standard method for generating the PSD, and is typically similar to what is used in, e.g., the firmware running on board of the OMB-v2018 and OMB-v2021 rabault2020open ; rabault2022openmetbuoy , and similar buoy designs kohout2015device . We also performed consistency checks slightly modifying these parameters during the exploration phase of this work, which confirmed that the results were not affected.
High-frequency peaks in the PSD of the wave motion are easier to visualize in the acceleration spectra rather than the elevation spectra. This is because, at the hardware level, the actual physical sensors used in the IMUs measure the acceleration, not the elevation, of the waves. Therefore, converting from the acceleration spectrum to the elevation spectrum requires applying the scaling $\mathrm{PSD}_{\eta}=\omega^{-4}\mathrm{PSD}_{\mathrm{acc}}$, where $\mathrm{PSD}_{\eta}$ is the PSD of the elevation and $\mathrm{PSD}_{\mathrm{acc}}$ is the PSD of the acceleration tucker2001waves . Thus, the spectral noise level goes from a constant (horizontal line) in $\mathrm{PSD}_{\mathrm{acc}}$ to a sloped line in $\mathrm{PSD}_{\eta}$, which makes it harder to visually distinguish noise from signal in the elevation spectrum. As a consequence, we consistently show acceleration spectra rather than elevation spectra in the following.
2.2.3 Cross Spectral Density and dispersion relation calculation
The cross spectral density (CSD) between two adjacent sensors can be used to recover the frequency-dependent phase shift due to the wave propagation delay between the two sensors and hence the dispersion relation longuet1963observation ; tucker2001waves ; marchenko2017field . This analysis can be performed on the T15 dataset; for other datasets, we do not have data from 2 or more adjacent sensors.
When performing a CSD analysis between two sensors (here, generally indicated by 111 and 222) to obtain the dispersion relation, the following procedure is used:
  • The cross-spectral coherence is calculated following the formula: $\mathrm{CSC}_{1,2}(f)=\frac{|P_{1,2}(f)|^{2}}{P_{1,1}(f)P_{2,2}(f)}$, where $P_{i,j}(f)=\mathrm{FFT}_{i}(f)\mathrm{FFT}_{j}(f)^{*}$ is the cross-spectral density between the IMUs number $i$ and $j$, $\mathrm{FFT}$ indicates the Fast Fourier Transform operator, and indicates complex conjugate. In practice, this is evaluated using the Welch method, i.e. averaging over sub-segments, similar to what is done for the PSD. A coherence value significantly above 0 indicates that phase locking is observed between the two sensors for the frequency considered, and that it is meaningful to extract a phase shift value from the cross-spectral density $P_{1,2}$. Here, we fix the threshold for the cross-spectral coherence at a value of 0.25. This is somewhat low, but this choice is dictated by the low signal-to-noise ratio in the acceleration data, as visible in the PSDs, so that the signal-to-noise ratio of the CSC is also relatively limited. In the frequency domain around the high-frequency peak, where the signal-to-noise ratio goes down to around 5, we even consider a lower coherence value threshold of 0.15, though we acknowledge that this is pushing the limit and the results should be considered with care and only viewed as an indication. However, as visible in Fig. 3, the results are consistent and correspond well to the theory, giving us a posteriori confidence in these relatively low threshold values.
  • The phase shift of the cross-spectral density between sensors 1 and 2, $\Delta\phi_{1,2}(f)=\mathrm{phase}(P_{1,2}(f))$, is calculated at the frequencies for which coherence about the threshold is observed. Since the IMUs used in the T15 dataset are generally aligned with the direction of the fjord, which constrains the direction of wave propagation, we calculate the real wavenumber component as: $k_{r}(f)=\Delta\phi_{1,2}(f)/D_{1,2}$, with $D_{1,2}$ the distance that separates the two sensors.
2.2.4 Bicoherence calculation
The (auto-, also known as self-) bicoherence mendel1991tutorial ; he2022rossby is used to estimate the degree of phase-locking between 3 frequencies $f_{1}$, $f_{2}$, and $f_{3}=f_{1}+f_{2}$ in a given timeseries signal. Here, we use the classical definition of the norm-2 bicoherence:
$b^{2}(f_{1},f_{2})=\frac{|\sum_{n}F_{n}(f_{1})F_{n}(f_{2})F_{n}^{*}(f_{1}+f_{2})|^{2}}{\sum_{n}|F_{n}(f_{1})F_{n}(f_{2})|^{2}~\sum_{n}|F_{n}(f_{1}+f_{2})|^{2}},$ (5)
where the index $n$ indicates the Welch segment considered, and $F_{n}$ is the FFT of the n-th segment.
Consider the limit case where there is perfect phase locking between components 1, 2, and 3. In this case, all terms in the numerator have the same phase, and the bicoherence will be equal to 1. In the opposite limit case, where there is no phase locking between components 1, 2, and 3, the numerator will average out to 0 given enough samples, and so will the bicoherence. Therefore, a bicoherence of 0 (or close to 0 for a finite and limited number of samples $n$) indicates that there is no phase locking between the components at $f_{1}$, $f_{2}$, and $f_{3}$. The higher the bicoherence value, the more phase locking, with a perfect phase-locking corresponding to a bicoherence of 1 (in practice, a value of 1 is never obtained, due to noise and uncertainties). Note that, as a consequence, a high value of the bicoherence is a necessary but not sufficient condition to demonstrate nonlinear interaction between the frequencies $f_{1}$, $f_{2}$, and $f_{3}$. The bicoherence contains additional information compared with the PSD per se: it is possible to generate signals with similar PSDs but very different bicoherence, as illustrated, for example, in the exercise available at https://github.com/jerabaul29/tutorials/blob/main/Bicoherence/Bicoherence.ipynb (accessed 2025-07). However, the bicoherence per se is not enough to, e.g., determine the existence, intensity, and direction of possible nonlinear energy transfers.
2.2.5 Triads diagram calculation
Wave triads, i.e. 3-wave resonant nonlinear interactions, are a possible way through which nonlinear phenomena can transfer energy between wave components. As discussed in the introduction, the necessary conditions for triads to have the possibility to exist are described by Eqn. (3). Similarly to the bicoherence, these 2 conditions are necessary, but not sufficient, for nonlinear wave triad interaction to occur.
The canonical way to visualize possible triad combinations is to plot a two-dimensional triads diagram. For this, the conditions in which both triad conditions are fulfilled are plotted with the longest wave vector ($\mathbb{k}_{3}$) starting at the origin and extending along the x-axis. For each position of $\mathbb{k}_{3}$ on the x-axis, the locus of admissible $\mathbb{k}_{1}$ is then plotted. Therefore, any triangle with vertices (i) the origin, (ii) a locus of admissible $\mathbb{k}_{1}$, and (iii) the corresponding $\mathbb{k}_{3}$, represents a wave triad condition. This is visible in the results in Fig. 4.
We generate the triads diagram for the T15 case in the following way. The dispersion relation that matches the most closely the one obtained from the cross-spectrum analysis is selected. Then, we select a set of values for $\mathbb{k}_{3}$. For each of these, the upper right quadrant of the two-dimensional plane is discretized numerically, with each point there corresponding to a candidate $\mathbb{k}_{1}$ and an associated candidate $\mathbb{k}_{2}=\mathbb{k}_{3}-\mathbb{k}_{1}$. For each of these, we then compute the residual $\mathrm{res}=f_{1}+f_{2}-f_{3}$, where $f_{1..3}$ is obtained by numerically solving the dispersion relation $D(k_{1..3},f_{1..3})$ for the corresponding point in the plane. Finally, the locus of $\mathrm{res}=0$, if it exists, is plotted as a level line. This indicates the locus of admissible $\mathbb{k}_{1}$ for which a triad ($\mathbb{k}_{1}$, $\mathbb{k}_{3}-\mathbb{k}_{1}$, $\mathbb{k}_{3}$) is obtained.
2.3 Analysis of data transmitted over Iridium
The OMB and Sofar Spotter datasets collected over Iridium satellite communications can transmit only a very limited amount of information, typically the 1-dimensional PSD of the wave motion with the OMB series of buoys, as discussed above. This makes it impossible to perform advanced timeseries analysis such as the bicoherence calculation described above.
Therefore, analyzing Iridium-transmitted buoy data is limited to identifying potential high-frequency peaks that may be associated to nonlinear wave interactions. High-frequency peaks in the PSD transmitted by autonomous buoys should be seen mostly as hints, not as a proof, that nonlinear interactions may take place, since it is in theory possible to have high-frequency peaks that do not correspond to a phase-locked mechanism. However, the existence of high-frequency peaks that are not related to the main peak in the PSD requires an arguably unlikely coincidence to take place, which is why we value documenting the recurring existence of such high-frequency peaks.
3 Results
3.1 Tempelfjorden 2015 (T15)
The T15 dataset contains clear evidence of high-frequency spectral peaks for some of the time segments considered. Fig. 1 presents an example of a segment with (left, segment 2) and without (right, segment 16) a high-frequency peak. The corresponding bicoherence plots are presented in Fig. 2. This clearly shows that a significant bicoherence peak is obtained in the case where the high-frequency spectral peaks is present. The frequencies associated with the bicoherence peak are typically located around $f_{1}\approx 0.16$, $f_{2}\approx 0.18$, resulting in $f_{3}=f_{1}+f_{2}\approx 0.34$, which is well matched with the high-frequency peak observed in the PSD. The bicoherence peak is not a sharp 1-point extremum but rather a "blob". This indicates that nonlinear interactions may take place over some ranges of frequencies and wavenumber, in good agreement with the observation of a relatively wide high-frequency peak in the PSD. We consider the sensitivity of the triad diagram to variations of these frequencies, among other parameters, later in this manuscript.
Note that the high-frequency peak is observed in the IMUs VN1002 and VN1007, which are farthest into the ice, but not in the IMU VN1005, which is closest to the ice edge. This may indicate either that the high-frequency peak is locally generated in the ice and does not exist close to the open water, or that the high-frequency energy distribution presents a nodes and antinodes spatial pattern reminiscent of a standing wave, or that some other complex variation of the spatial distribution of the high-frequency energy content is present in this case.
Figure 1: PSDs obtained from the 3 VN100 IMUs that are part of the T15 dataset. Left: PSDs from the segment number 2. A clear high-frequency peak is observed in IMUs VN1002 and VN1007 (farther in the ice), but not in VN1005 (closest to the open water). The high-frequency peak is relatively broad. Right: PSDs from the segment number 16. No high-frequency peak is observed.
Figure 2: Bicoherence obtained from the IMU VN1002 from the T15 dataset. Left: bicoherence from the segment number 2. A clear peak in the bicoherence is observed, typically around the frequencies $f_{1}\approx 0.16$, $f_{2}\approx 0.18$, resulting in $f_{3}=f_{1}+f_{2}\approx 0.34$, which matches the main and high-frequency peaks in Fig. 1. Note however that the bicoherence peak is rather a "blob" than a one-point extremum, in good agreement with the observation of a relatively broad high-frequency peak in the PSD plot. Right: bicoherence from the segment number 16. No peak is observed in the bicoherence.
In order to further investigate whether the high-frequency spectral peak and bicoherence peak can correspond to a nonlinear wave interaction caused by wave triads, we compute the real-world wave dispersion relation by applying the cross-spectral density analysis between the VN1005 and VN1007 IMU sensors. These are the 2 IMUs that are located furthest away from each other, which gives a higher relative accuracy in the phase-shift measurements. Due to the separation distance, phase aliasing is observed. This phase aliasing is removed by enforcing a smooth monotonic increase in $\Delta\phi$ for increasing frequencies, by applying suitable corrections by $2\pi$ when aliasing jumps take place. The results are presented in Fig. 3, with the dispersion relation from Eqn. (3) included for different effective sea ice thickness values and otherwise standard landfast ice mechanical characteristics (we use similar values as rabault2016measurements , i.e. $g=9.81$ m/s2, $E=2\times 10^{9}$ Pa, $\nu=0.3$, $\rho_{w}=1025$ kg/m3, $\rho_{i}=920$ kg/m3). As visible in Fig. 3, and in good agreement with the findings previously reported in sutherland2016observations , a significant deviation from the open water dispersion relation, which matches well with a thin-plate elastic dispersion relation with the same typical thickness as measured in the field, is observed early in the data (up to around segment 4). We show the data points obtained about the low signal-to-noise ratio higher frequency components by using the low CSC threshold as thin crosses, and we expect that these are particularly noisy and should, therefore, only be taken as an indication. Later on, the dispersion relation collapses back to the deep-water dispersion relation, which we interpret to correspond to the development of cracks through the ice cover. Cracks were indeed observed when the instruments were recovered, while they were not present when the instruments were deployed, consistent with this interpretation.
Figure 3: Empirical dispersion relation obtained from a cross-spectral analysis between the IMUs VN1005 and VN1007. Clear deviation from the deep water dispersion relation, in good agreement with the thin elastic plate dispersion relation and with previous results sutherland2016observations , is observed until around segment 5. For later segments, the empirical dispersion relation matches the deep water one. Here, we show both the empirical dispersion relation obtained for the main PSD peak (large dots), and for the higher frequency components when these are present in the PSD (crosses; due to the low signal-to-noise ratio in the higher frequencies, these are obtained using a low value of the CSC threshold, and are therefore particularly noisy and should be taken only as an indication). The best-matching dispersion relation, which corresponds to an ice thickness of around 0.6 meters with the default ice parameters listed in the text (in good agreement with the field measurements reported in sutherland2016observations ), provides the $B_{\mathrm{ref}}$ value for the $B$ coefficient in Eqn. (2).
We use the best-match elastic plate dispersion relation from Fig. 3 (i.e., the one for an elastic plate with thickness 0.6m) to perform a wave triad analysis. The corresponding value for the $B$ coefficient in Eqn. (2), is marked as $B_{\mathrm{ref}}$. The results are presented in the wave triad diagram in Fig. 4. As visible there, triads are indeed possible given this dispersion relation. The specific locus of $\mathbb{k}_{1}$ corresponding to the value of $k_{3}$ that matches the frequency for the high-frequency peak revealed by the bicoherence analysis is plotted in black. For this locus, we find two sets of wave vectors $(\mathbb{k}_{1},\mathbb{k}_{2})$ that match the frequencies $(f_{1},f_{2})$ (resp., $(f_{2},f_{1})$) obtained from the bicoherence peak. These are plotted in red. Therefore, the PSD, bicoherence analysis, dispersion relation derived from the CSD analysis, and wave triads diagram are all consistent with each other and indicate that nonlinear mechanisms corresponding to wave triads are compatible with our observations. Note that all these empirical facts are necessary and not sufficient conditions to establish the existence of wave triads. Therefore, this is the strongest set of evidence we can provide with the dataset at our disposal, but this is not an indisputable proof that our observations arise due to nonlinear wave triads redistributing energy across spectral components.
Knowing that triads can be expected, one can also observe, back on Fig. 1, that there are two significant dips in the main peak of the PSD of the IMUs furthest in (VN1002 and VN1007) relatively to the PSD of the IMU furthest out (VN1005). These dips are observed consistently over several FFT points, for both IMUs further in the ice, at frequencies around 0.150.150.15Hz and 0.190.190.19Hz, which sum up to the typical frequency of the high-frequency peak (around 0.340.340.34Hz). These may be a signature of the energy that is being "stolen" from the main peak and transferred into the high-frequency peak by the triad interactions.
Figure 4: Wave triad diagram corresponding to the thin plate dispersion relation that best matches the empirical dispersion relation observed for early segments in Fig. 3 (i.e. elastic plate with thickness 0.6m). This dispersion relation can sustain the wave triad conditions in the frequency range of the high-frequency peak that we observe. For the specific value of $\mathbb{k}_{3}$ corresponding to the PSD and bicoherence high-frequency peak, sets of locii for $\mathbb{k}_{1}$ and $\mathbb{k}_{2}$ that match the main peak in the PSD are identified (red). In this and similar figures, $kx$ and $ky$ indicate the $x$ and $y$ components of a wave vector $\mathbb{k}=kx\mathbb{e}_{x}+ky\mathbb{e}_{y}$.
Since these data are obtained from field measurements, there are significant uncertainties on all parameters related to the ice properties (which are known to vary widely in time and space). In addition, the bicoherence peak is a relatively large "blob", and the PSD high-frequency peak is also relatively wide. Therefore, there is no single clear value for these parameters. Moreover, the "peak bicoherence triad" analyzed above corresponds only to the maximum of the bicoherence "blob". In reality, the spectra observed are broad-banded, and the high-frequency peak may rather be the consequence of a continuum of triad interactions than of a single triad. In particular, a full description of the dynamics would require integrating over all possible triads, akin to Hasselmann’s equation hasselmann1962non . However, this is beyond the scope of the present work, and to offer a convincing comparison between these more advanced descriptions and observations, one may need significantly more detailed field data including measurements from a large array of instruments.
In order to double-check the robustness of our analysis, we therefore perform a sensitivity analysis by varying the triad frequencies and the $B$ coefficient in Eqn. (2). The results, which cover the typical range of admissible parameter values, are presented in Fig. 5. As visible there, our results are robust to variations of the $f_{1}$, $f_{2}$, and $f_{3}=f_{1}+f_{3}$ frequencies within the range of values observed within the bicoherence blob, even for quite large changes in the coefficient $B$. This gives us good confidence in the robustness and validity of our results.
Figure 5: Sensitivity analysis of the triads shapes. The 3 plots indicate the triads found for different values of $f_{1}$, $f_{2}$, and $f_{3}=f_{1}+f_{2}$ chosen to span the width of the blob of high bicoherence observed in Fig. 2. The series of triad diagrams for each of the plots, coded by the colors, corresponds to different values of $B$ relative to the best fit value $B_{ref}$ found in Fig. 3. Each triad locii curve is for the value of $k_{3}$ obtained from solving the dispersion relation for $f_{3}$ using the corresponding value of $B$. This confirms that our results are quite robust to varying all the parameters involved.
3.2 Yermak plateau 2020 (Y20)
Similarly to the T15 dataset, the Y20 dataset contains clear high-frequency peaks in the PSDs of some data segments, as visible in Fig. 6, which compares one case without (left, segment 317) and with (right, segment 322) such a high-frequency peak. The associated bicoherence for the same 2 segments is presented in Fig. 7. A clear bicoherence peak, which corresponds to the high-frequency peak observed in Fig. 6, is visible there.
Since only one buoy was deployed, the dispersion relation cannot be recovered from the CSD as was done in the T15 case. Therefore, we have no way to know with certainty how the dispersion relation exactly looks like in this case and what the associated triads diagram is. However, there is no doubt that, similar to the T15 case, a dispersion relation that contains a significant elastic plate contribution will generate triads that can be compatible with the bicoherence observed.
Figure 6: PSD of the vertical acceleration recorded by the VN100 IMU present in the Y20 buoy. Left: example of PSD without high-frequency peak present (corresponding to segment 317). Right: example of PSD with a high-frequency peak (corresponding to segment 322).
Figure 7: Bicoherence of the vertical acceleration recorded by the VN100 IMU present in the Y20 buoy. Left: bicoherence for the same segment that does not contain a high-frequency peak as used in Fig. 6 (segment 317). Right: bicoherence for the same segment that does contain a high-frequency peak as used in Fig. 6 (segment 322). A clear bicoherence peak, that matches the main PSD peak and the high-frequency peak, are observed for segment 322.
3.3 Spectra over Iridium
Most available waves in ice data rely on Iridium communications for data collection and report only the 1-dimensional PSD of the wave motion. This makes it impossible to compute the bicoherence. However, it is still possible to identify high-frequency peaks in PSDs, which may indicate candidate events for dynamics similar to what has been identified in the T15 and Y20 datasets.
Some wave events with clear high-frequency peaks, observed in Antarctica, are presented in Fig. 8. These data were collected by 3 different kinds of instruments: OMB-v2021 (first row), OMB-v2018 (second row), and Sofar Spotter (third row). Clear second-order high-frequency peaks are observed, and in some cases even third-order high-frequency peaks are visible.
Figure 8: Examples of high-frequency peaks in the PSD of waves in ice data from Antarctica, from the Casey 2022 deployment (using OMB-v2021, first row) and the Davis 2020 deployment (using OMB-v2018, second row, and using Sofar Spotter, third row). Clear high-frequency peaks are visible, across all 3 instrument kinds, sometimes even with a third-order peak being present.
Similarly, a few similar events collected by OMB-v2021 in the Arctic are presented in Fig. 9. There, too, cases with both second- and sometimes third-order high-frequency peaks are visible. Note that the AWI-UTOKYO data show the PSDs reported by 3 different buoys over 3 different time periods, so these cover several events. More events (not all reproduced here) are visible in the data. In addition to the presence of the high-frequency peaks by themselves, a number of interesting features are visible:
  • The strength of the high-frequency peak in the CIRFA-UIT 2022-07 plot seems to be very sensitive to the exact features of the main peak. In particular, a much higher high-frequency peak is observed for the light-green data segment, while this corresponds only to a subtle change (slightly higher peak frequency) in the main peak. This may be seen as a possible signature of strongly nonlinear mechanisms being at play.
  • In CIRFA-UIT 2022-08, the peak frequency of the main high-frequency peak is slowly increasing over time. This is associated with a well-matched shift of the frequency of the high-frequency peak to higher frequencies. This seems to confirm the tight coupling between the main peak and the high-frequency peak, i.e., that these two are coupled and change in unison, and are not just unrelated peaks coming from unrelated phenomena that are here because of a coincidence.
  • In AWI-UTOKYO 2022-07, it seems that the presence of the high-frequency peak can be very sensitive to the exact shape of the main peak.
  • In AWI-UTOKYO 2022-08, the shape of the main peak remains quite similar, with a moderate reduction in amplitude occurring. However, the relative reduction in the amplitude is much greater in the high-frequency peak. Although the main peak is reduced by a factor of around O(10) between the first and last spectra (from around $4e^{-2}$ to around $5e^{-3}$), the high-frequency peak is reduced by a factor of around O(100) (from around $2e^{-4}$ to around $3e^{-6}$), in a way that is compatible with a second-order effect. A similar effect is observed in (f).
Figure 9: Examples of high-frequency peaks in the PSD of waves in ice data from the Arctic. Second, and sometimes even third order high-frequency peaks, are observed. The examples cover a range of deployment dates, locations, ice conditions, and use both OMB-v2018 and OMB-v2021. Several features that can be interpreted as possible consequences of nonlinear dynamics are visible, as described in the text.
In addition to these observations, we also considered the PSD spectra from deployments in the MIZ in the case where small broken ice floes were encountered. These include in particular the outermost buoys in the dataset 2021-03 of rabault2023dataset , see more descriptions in nilsen2021pc , and the outermost buoys in the datasets of muller2025distributed ; muller2025distributeddata . There, we could not find clear high-frequency peaks similar to those reported in Fig. 9. This further hints at the importance of having an ice sheet or a large enough ice floe to support significant flexural effects for observing the high-frequency peaks we report.
Since the bicoherence analysis cannot be conducted based on the 1-dimensional PSDs that are available to us alone, and we do not have enough information to confidently determine the dispersion relation and associated wave triad diagrams, we cannot assert with certainty that these high-frequency peaks are phase-locked with the main peak nor that they are compatible with the local dispersion relation to correspond to wave triads. Therefore, these observations should only be considered as circumstantial evidence for the possible existence of nonlinear mechanisms such as wave triads. However, the relatively common occurrence of these high-frequency peaks across a range of conditions, deployments, and instruments confirms that they most likely correspond to a real physical phenomenon and are not a seldom one-time coincidence or technical artifact.
4 Discussion
We have considered 3 kinds of in-situ observation dataset: (i) the T15 dataset, where the timeseries from several closely-located time-synchronized IMUs deployed in Arctic landfast ice are available, (ii) the Y20 dataset, where the timeseries from a single IMU-based buoy deployed in the Arctic was recovered, and (iii) a number of deployment datasets in both the Antarctic and the Arctic, where only 1-dimensional PSDs are recovered over Iridium satellite communications.
In all cases, clear high-frequency peaks can be observed in some of the data segments. Since these high-frequency peaks are clearly above the noise background, are sometimes present and sometimes not (so they are not a systematic noise in the sensors and / or processing algorithms), and these can be observed over a range of instruments (VN100-based loggers, OMB-v2018 with a VN100 IMU, OMB-v2021 with an ISM330DHCX motion sensor, Sofar Spotter using GPS Doppler data) and deployments in various sea ice conditions and locations covering both the Arctic and the Antarctic, we can be confident that these are not noise nor a one-off coincidence, but a real recurring physical signal.
In the cases where the timeseries are obtained (T15 and Y20), a bicoherence analysis confirms that the maximum of the high-frequency peak is phase-locked with 2 components of the main peak. This is a necessary (but not sufficient) condition for the high-frequency peak to be the product of e.g. nonlinear wave triad interactions or some other nonlinear physical mechanism related to the main peak in the PSD.
In the one case where timeseries from several closely located and time-synchronized sensors are obtained (T15), the dispersion relation is recovered from a CSD phase-shift analysis. The results indicate that in the signal segments for which the high-frequency peak is observed in the PSD and is visible in the bicoherence analysis, clear flexural-elastic thin plate effects are present in the dispersion relation, which is compatible with the formulation of Eqn. (2). This elastic-flexual component makes it possible for the dispersion relation for the waves propagating in the ice to support wave triads. We show that the associated wave triads diagram indicates that matching triads can exist between the frequencies revealed by the bicoherence analysis.
A critical question to elucidate regarding the high-frequency spectral peaks is whether these correspond to bound or free waves. Bound waves, which canonical example is the higher-order components of the Stokes solution lamb1924hydrodynamics , are perturbations to the first-order wave solution that are present as the higher-order consequence of nonlinear terms. These can arise either from the effect of nonlinear terms on a single wave mode (as is the case with the higher order Stokes wave components), or as the effect of nonlinear terms on a superposition of wave modes (as the presence of nonlinear terms implies that the linear superposition of two or more wave modes results in an additional nonlinear residual term). Bound waves do not follow the dispersion relation, and are caused by and travel together with the main propagating wave, materializing as "disturbances" of the main wave. Free waves, on the contrary, propagate freely and follow the dispersion relation. Based on Fig. 3, it seems most likely that the high-frequency peak observed in the T15 dataset corresponds to a free wave that follows the dispersion relation. However, we acknowledge that the signal-to-noise ratio used there is low, so offering a definite confirmation that our high-frequency spectral peaks correspond to free waves would require additional observation and confirmation.
These results can be seen as a strong hint that nonlinear dynamics are a likely cause of the high-frequency peaks observed and that triads are a strong candidate for the underlying energy-transfer mechanism. However, we want to highlight that our results are proofs that the necessary conditions for triads are fulfilled, but that these are not strictly speaking sufficient conditions / observations to assert with complete certainty that triads are indeed the underlying mechanism. To bring certainty to this matter, additional measurements designed specifically to test this hypothesis would be needed, as suggested below. One could also study the series expansion of the nonlinear water wave equations under conditions that correspond to the flexural-gravity wave dispersion relation, and estimate the nonlinear interactions that can arise from these, or perform fully nonlinear simulations.
An additional observation is that wave directionality seems to play a role. While the frequency and wavenumber triad conditions provide necessary conditions for triads to be possible, these say nothing about the strength and direction of the energy transfers between the 3 frequencies involved. In particular, the extraction of energy from the two low-frequency components and transfer to the high-frequency one may only happen for some particular frequencies and relative orientation conditions. For example, in the case of the T15 dataset, we note that the candidate triad occurs for wave components 1 and 2 that have a relatively large relative angle, keeping in mind that this is observed deep in a narrow fjord system (and even changing a bit the parameters involved, significant angle is observed, as long as one does not go to the extremes of the parameters that can sustain triads). Similarly, a quick investigation of when high-frequency peaks are obtained in Antarctica seems to indicate that these happen mostly relatively close to open water (this is based on a visualization of the sea ice concentration maps in the area, not reproduced here). This may be consistent with the fact that, as waves propagate through the ice, the directional spectrum becomes more narrow (since components going sideways have longer effective propagation distances and are, therefore, more strongly attenuated). However, we consider that an exhaustive and detailed meta-studies of when and where these high-frequency peaks occur in all OMB and other wave in ice datasets openly available, and a comparison to incoming wave spectra and sea ice conditions, would be a major effort that is out-of-scope for the present manuscript.
While our observations seem to be compatible with the existence on nonlinear wave interaction and energy transfers between frequency modes through a wave triad mechanism, other mechanisms could also play a role and / or explain our observations. For example, zhao2017gap suggests that arrays of adjacent floating bodies (in our case, ice floes) can lead to nonlinear resonance and create high-frequency peaks in PSDs. However, this seems unlikely to be the cause of the observations in T15. Indeed, in this case, the high-frequency peak is observed when a flexural dispersion relation is present, i.e. when the ice is unbroken, while it disappears when the flexural component disappears, i.e. the floes get broken. This is the opposite of what could be expected from the floes array resonance mechanism discussed in zhao2017gap . Other possible mechanisms could involve, for example, edge waves, which can also participate in sustaining nonlinear resonant energy transfers marchenko1999parametric . Note that, to make the matter even more complex, nonlinear energy transfers through resonant triads could even happen across different kinds of waves phillips1981wave (edge waves being a possible candidate). There may also be other mechanisms that we have not yet imagined and that could also result in these high-frequency peaks.
In order to answer these questions with certainty, a purposely-designed set of experiments can be envisioned.
  • Ideally, in the field, one should perform detailed time-resolved and time-synchronized wave acceleration measurements in a wide area of sea ice, in order to be able to resolve the full dynamics of the two-dimensional ice sheet. This is a realistic attempt anno 2025, thanks to the low cost and accuracy of modern motion sensors such as the ISM330DHCX used in the OMB-v2021 (cost O(20 USD)), and the time accuracy provided by GPS Pulse Per Second (PPS) information (UTC second start can be determined with a jitter of at most O(10 ns) using GPS chips that cost O(30 USD). This was used, for example, in voermans2023estimating , which designed and deployed a low-cost open source logger that could reach microsecond accuracy in signal acquisition timestamps and logging rates of several kHz on a SD card, using a low-cost micro controller. With proper engineering, an open source SD-based wave in ice logger could be built with high-accuracy wave measurements and (sub-) microsecond timing accuracy for O(150 USD) including all material costs. This means that the hardware for a large-scale measurement campaign with 200 such loggers deployed as an array on the sea ice would cost O(30k USD), well within reach of a small- to medium-sized scientific project. This would enable a detailed analysis of the in-situ dispersion relation, spatial and spectral distributions of the energy, and of the effect of directionality.
  • In addition, detailed analytical calculations, as well as numerical simulations of the waves in ice equations using a 3-dimensional nonlinear solver and considering the effect of the full directional spectrum, could help to understand the energy transfers at stake constantin2018nonlinear .
  • Finally, it could be possible to investigate these phenomena in a laboratory experiment. However, this may be challenging due to (i) the need to properly scale the physical phenomena happening, which can result in several non-dimensional quantities conflicting with each others as it has previously been observed in the study of waves in ice attenuation rabault2019experiments , (ii) the likely need to consider fully 2-dimensional directional wave spectra, (iii) inherent challenges related to reflections and standing modes in a wave tank, especially when the effects to measure are subtle higher order effects that can easily be hidden by larger systematic errors induced by wave reflections, imperfect paddle-matching conditions, and finite-length effects of either the wavetank or the wave packets generated, as highlighter previously e.g. in rabault2016ptv ; marchenko2021laboratory .
Before some additional, purposely gathered data are available, a natural extension to the present work is to perform an exhaustive meta-study on all available OMB and similar waves in ice data. In particular, it would be interesting to systematically compare when high-frequency peaks are observed with both the position of the buoys in the MIZ, the incoming wave conditions, the atmospheric conditions, and satellite imaging of the local sea ice configuration, to attempt to reveal critical factors that cause these high-frequency peaks to be present or absent. We consider that such a study would require a major effort to be as representative as possible and to consider as many datasets as possible, and that it is beyond the scope of the present manuscript.
An interesting question is the impact and applicability of the present findings. While the present analysis gives us confidence that the spectral high-frequency peaks are real and may likely be related to interesting nonlinear dynamics, are these a physical and mathematical curiosity without much practical value, or are they important for sea ice dynamics or can be otherwise useful? If these high-frequency peaks could be reliably observed from satellite data, then they could possibly give information about the existence of flexural stresses in the ice and, hence, about the effective properties of the sea ice (in particular, through the information carried about the value of the $B$ factor in Eq. (2)). However, it remains to be seen whether these high-frequency peaks can be visible in satellite images and how much information can be inferred back.
Another interesting question is to determine how much additional energy dissipation can be induced by the nonlinear energy transfer to the high-frequency peak and the dissipation there. To estimate the importance of this effect, we can attempt to perform some order-of-magnitude calculations. To do so, we assume the following.
  • We work with a time-stationarity approximation: i.e., we assume that, averaged over a time O(20 minutes) corresponding to the duration over which the spectra presented above are computed, the spectra can be taken as a representative state of the wave energy distribution.
  • We neglect variability in space, and consider that the high-frequency spectral peaks observed are in a state of local energy balance (effectively neglecting gradients in the wave energy fluxes).
  • We consider that 2 mechanisms are involved in sustaining the high-frequency peak and determining its level: i) a net nonlinear energy transfer $S_{\mathrm{nl}}$ from the main peak to the high-frequency peak, and ii) wave energy dissipation $S_{\mathrm{diss},\mathrm{high-freq}}$ associated with the presence of the high-frequency peak.
Under these assumptions, the nonlinear energy transfer to the high-frequency peak and the dissipation there are in balance: $S_{\mathrm{nl}}=S_{\mathrm{diss},\mathrm{high-freq}}$. Therefore, estimating $S_{\mathrm{diss},\mathrm{high-freq}}$ provides an estimate for $S_{\mathrm{nl}}$. As a consequence, estimating the relative importance between $S_{\mathrm{diss},\mathrm{main-peak}}$ (the energy dissipation at the main peak) and $S_{\mathrm{diss},\mathrm{high-freq}}$ provides an indication of the fraction of the wave in ice attenuation in the main peak that comes from the nonlinear energy transfer and dissipation at higher frequency. To estimate this relative amount of energy dissipation between the main and the higher-frequency peaks, we use the following hypothesis.
  • We assume that the wave energy spatial dissipation rate is similar to what is derived by yu2022 , in particular their Fig. 4.b which is based on a large set of test data: $k_{i}\propto f^{4.5}$.
  • The temporal dissipation rate at a fixed location can be calculated as: $\beta=c_{g}k_{i}$ gaster1962note ; sutherland2017attenuation ; with deep water $c_{g}\propto f^{-1}$, this leads to $\beta\propto f^{3.5}$. Therefore, at a fixed location, the temporal dissipation rate ratio between the energy dissipation in the high-frequency peak and the main peak is $\mathrm{DR}_{2f}/\mathrm{DR}_{f}=2^{3.5}\mathrm{PSD}_{\eta,2f}/\mathrm{PSD}_{\eta,f}$, where $\mathrm{DR}_{f}$ is the temporal dissipation rate at frequency $f$.
  • $\mathrm{PSD}_{\eta}\propto f^{-4}\mathrm{PSD}_{\mathrm{acc}}$, and we are looking at $\mathrm{PSD}_{\mathrm{acc}}$ in our graphs, while the scaling formula for $k_{i}$ applies to $\mathrm{PSD}_{\eta}$; therefore, $\mathrm{DR}_{2f}/\mathrm{DR}_{f}=(2^{3.5}/2^{4})\mathrm{PSD}_{\mathrm{acc},2f}/\mathrm{PSD}_{\mathrm{acc}f}\approx 0.7\mathrm{PSD}_{\mathrm{acc},2f}/\mathrm{PSD}_{\mathrm{acc},f}=0.7R_{f,2f}$, where $R_{f,2f}=\mathrm{PSD}_{acc,2f}/\mathrm{PSD}_{\mathrm{acc},f}$
Taking into account the PSD plots presented in the figures above, $R_{f,2f}$ shows quite a large degree of variability. For some of the PSDs in the Antarctic $R_{f,2f}\approx O(1)$ ("A1"); for T15 $R_{f,2f}\approx O(0.1)$; for most other cases, we generally observe $R_{f,2f}\in{[O(0.1),O(0.01)]}$ ("O"). Therefore, our order-of-magnitude estimates based on the stationarity approximation indicate that damping in the high-frequency peak may represent up to 70% of the intrinsic energy dissipation for the main peak in the case "A1", 7% in the case T15, and 7% to 0.7% in the "O" cases. This indicates that this mechanism may, in certain circumstances, play a significant role (nearly doubling the amount of energy dissipation in the case "A1"), but that it may also in many cases be only a minor contribution to the total wave energy dissipation. However, note that these are only very coarse order-of-magnitude estimates. For example, we do not take into account the spectral width of the main vs. high-frequency peaks. Moreover, the high-frequency peak dissipation effect may already be somehow present in the wave in ice empirical dissipation laws that are fitted to observations, in which case the exponent in the frequency dependence of the energy dissipation may be biased. Therefore, additional and more detailed investigation is needed to provide a definitive conclusion, and one may need to consider a systematic, representative, large-scale, and well-sampled dataset to cover all typical cases of ice and wave conditions.
5 Conclusion
We show that high-frequency spectral peaks can be reliably observed in the 1-dimensional Power Spectrum Density (PSD) of waves in ice buoys. Although these are only intermittently observed, we show that they occasionally occur in a wide range of deployments sampling various sea ice conditions using several types of instruments. In the cases where in-situ timeseries and not just 1-dimensional PSDs are obtained, we show that these high-frequency peaks are associated with a high value of the (self-) bicoherence. In a specific case where several sensors are placed adjacent to each other on the sea ice, we show that the empirical dispersion relation obtained from a cross-spectrum density (CSD) analysis contains significant flexural effects, which makes it compatible with the existence of wave triads that match the frequencies of the high-frequency peak and the bicoherence peak.
Our results indicate that a strong set of concomitant evidences, which are compatible with each other and mutually reinforcing, suggests that waves in ice high-frequency spectral energy peaks are a relatively common mechanism and may be associated with wave triad interactions. This, in turn, can extract energy from the main spectral energy peak, and re-inject this energy into a higher-frequency, more heavily damped peak. As a consequence, this mechanism can introduce a significant contribution to the wave in ice dissipation.
Although it seems very unlikely for our observations to be the result of a coincidence, the results that can be derived from our data are all necessary but strictly speaking not sufficient conditions to establish that wave triad interaction is truly the underlying mechanism causing these high-frequency peaks. Therefore, further investigation is necessary to definitively confirm that these high-frequency peaks are induced by wave triads that cause nonlinear energy transfer from the main PSD peak to the high-frequency peak(s), and quantify the energy transfers in detail. We suggest that such further investigations can be based on a combination of more detailed and advanced field measurements, numerical studies, and laboratory experiments.
6 Acknowledgments
The help of the UNIS logistics team during the 2015 T15 fieldwork is gratefully acknowledged. The help of laboratory engineer Olav Gundersen (UiO) when building and preparing the equipment used in the T15 and Y20 deployments is gratefully acknowledged. We would like to thank the crew of the RV Kronprins Haakon for invaluable help deploying the buoys in the Arctic. The authors thank Pr. Yngve Kristoffersen (UiB) for having deployed the buoy used in the Y20 dataset. Funding from the Research Council of Norway through the projects "Waves in Oil and Ice" (WOICE, project 233901), "Dynamics Of Floating Ice" (DOFI, project 280625), "Arven Etter Nansen" (AeN, project 276730), "FOCUS" (project 301450), is gratefully acknowledged.
7 Use of Artificial Intelligence tools in the writing of this manuscript
Artificial intelligence tools, in particular Large Language Models (LLMs), were used for the writing of this manuscript. The scientific content of the work was the product of the work of the authors, and the LLMs were used only to help polish the language and the formulation of the prose. In particular, the authors reviewed and quality controlled each contribution from the LLMs.
Appendix A: open data and open code
The data used in this manuscript are openly available.
Some of the code used in this manuscript is openly available at the following URL: [will be made available upon peer-reviewed publication of the manuscript]. We invite readers who have comments or are willing to interact with us to do so through the Github issue tracker of the corresponding repository.